OpenStudy (anonymous):

A car starts at 10.00 am at speed of 60km/h and 2nd starts one hour later at 80km/h. at what time will the 2nd car overtake the 1st?

OpenStudy (anonymous):

Key: Distance covered by both the cars is same. So, if the 2nd car overtakes the 1st car after x hours (from 10:AM), then Distance covered by the 1st car = 60 times x km and that covered by 2nd car = 80 time (x-1) so, 60x = 80x- 80 x= 4 hours At 10 + 4 = 12 Overtake will be at 12 noon

OpenStudy (anonymous):

10 + 4 = 2 pm thnksss

OpenStudy (anonymous):

Oh..! :P

OpenStudy (anonymous):

yea :P